Rory, Leo, and Joey win $362 880 in a lottery. They decide to divide the winnings so
that Rory gets two parts, Leo three parts, and Joey four parts. Exactly how much money
does each person receive?

Answer by addingup(3677)

2 parts + 3 parts + 4 parts = 9 parts. How many 9 parts are in 362880?
There are:
362880/9 = 40,320
Now distribute:
40,320*2 = 80,640 This is Roy
40,320*3 = 120,960 This is Leo
40,320*4 = 161,280 This is Joey
Total . .= 362,880 Correct